this year (10 years after i took out my first loan) i check my loan balance. Only part of my payments have been going to pay down the loan; the rest has been going towards interest. I see that i still have $88,536 left to pay on my loan. My house is now valued at $160,000. How much of the original loan have i paid off?
1
Expert's answer
2020-07-16T16:45:13-0400

S = P + I and I = Prt

S= P +Prt

S =simple interest

P= present value

I= interest earned

r= rate of interest

t= time in years

S = P(1+rt)

given , P= $160,000 and t = 10 year

r = ?

we find r through r=payment fund * payment period left /P

Payment fund = $88,536

Payment period left= 1 installment left

P=$160,000

r=88,536*1/160,000

r=0.55335%

now S = P(1+rt)

S=160,000{1+(0.55335)*10)}

S=160,000(1+5.5335)

S= 160,000(6.5335)

S=$1,045,360